OpenStudy (anonymous):

I need to take half the coefficient term of the x term and square the result. f(x)=1/3x^2+x+1

OpenStudy (anonymous):

\[f(x)=\frac{ 1 }{ 3 }x^2+x+1\]

OpenStudy (anonymous):

Simple, but I feel like I'm doing it wrong. Please explain.

OpenStudy (anonymous):

@rajathsbhat @kery

OpenStudy (anonymous):

ultimately what i need to do is write the given expression in form f(x)=a(x-h)^2+k, but i believe this is the first step

OpenStudy (anonymous):

^That clears things up.

OpenStudy (anonymous):

yes its about identifying the vertex and all the examples are shown without fractions

OpenStudy (anonymous):

Ok first write F(x) as \[\frac{1}{3}(x^{2}+3x+3)\]

OpenStudy (anonymous):

Now does it look familiar?

OpenStudy (anonymous):

i dont understand where is half the coefficient, then its square

OpenStudy (anonymous):

thats what it says to do first

OpenStudy (anonymous):

and then to add and subtract that value

OpenStudy (anonymous):

i dont want to get to far ahead but after that it says to factor the first three terms aas perfect square trinomial, those are the steps in order

OpenStudy (anonymous):

Yes. keep that 1/3 outside & do everything you just said to \((x^{2}+3x+3)\)

OpenStudy (anonymous):

i just dont understand how you got that

OpenStudy (anonymous):

would be final but i just cant get it \[f(x)=\frac{ 1 }{3}(x+\frac{ 3 }{ 2 })^2+\frac{ 1 }{4 }\]

OpenStudy (anonymous):

\[\large \begin{align}F(x)&=\frac{1}{3}(x^{2}+3x+3)\\&=\frac{1}{3}[x^{2}+3x+3+(\frac{3}{2})^{2}-(\frac{3}{2})^{2}]\\&=\frac{1}{3}[(x^{2}+3x+(\frac{3}{2})^{2})+3-(\frac{3}{2})^{2}]\\&=\frac{1}{3}[(x+\frac{3}{2})^{2}+\frac{3}{4}]\\&=\frac{1}{3}(x+\frac{3}{2})^{2}+\frac{1}{3}(\frac{3}{4})\\&=\frac{ 1 }{3}(x+\frac{ 3 }{ 2 })^2+\frac{ 1 }{4 }\end{align}\]

OpenStudy (anonymous):

Do you see what i did?

OpenStudy (anonymous):

so \[\frac{ 3 }{ 2 }\] is the result of half of the coefficient, and then squared?

OpenStudy (anonymous):

i thought half of 1/3 would be 1/6 and the square of 1/6 would be 1/36 - what am i doing wrong there

OpenStudy (anonymous):

1/3 is not the co-efficient of x for you to do that.

OpenStudy (anonymous):

so what is? i thought coefficient was value associated with x

OpenStudy (anonymous):

the way you should do these problems is first separate out the co-efficient of x^2. You see, X^2 should not have anything as its co-efficient. That screen-clip is a simple example that's all.

OpenStudy (anonymous):

is there anyway you can give a word description for those steps you did? simple terms would be ok i just have trouble recognizing what you did from step to step

OpenStudy (anonymous):

sure. I'll do that.

OpenStudy (anonymous):

1. I take 1/3 common from every term & put it outside the whole equation. Do you get it?

OpenStudy (anonymous):

Yes i think i am getting it thank you for your help. I will practice another question or two.

OpenStudy (anonymous):

yes. you do that. Good luck :)